# Landing with the Score: Riemannian Optimization through Denoising

### v1

@@ Appendix F Discretized Riemannian gradient flow and descent
$\sup_{x\in\mathcal{T}(\tau)}\lVert s^{\prime}(x)\rVert_{\infty}\leq\epsilon+\frac{1}{1-\tau/\tau_{\mathcal{M}}}\,.$
@@
$\lVert G_{s}^{\eta}(x)\rVert\leq C\lVert s^{\prime}(x)\rVert+\eta\lVert s(x)-x\rVert\leq C(\epsilon+\frac{1}{1-\tau/\tau_{\mathcal{M}}})+2\eta\tau\text{\ \ for\ \ }x\in\mathcal{T}(\tau)$
@@
F.1 Discretized approximate Riemannian gradient flow
@@
F.2 Proof of Theorem 5
Let us write $\epsilon^{\prime}=\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3}\leq\tau/2$. First we note that $\{x_{k}\}_{k=0}^{\infty}\subseteq\mathcal{T}(\epsilon^{\prime})$ as soon as $x_{0}\in\mathcal{T}(\tau/2)$, because $s(\mathcal{T}(\tau))\subseteq\mathcal{T}(\epsilon^{\prime})$ and $x_{k}-\gamma_{k}s^{\prime}(x_{k})\nabla f(x_{k})\in\mathcal{T}(\tau)$, since
$\gamma_{k}\leq\frac{\tau}{2C(\epsilon^{\prime}+(1-\tau/\tau_{\mathcal{M}})^{-1})}\leq\frac{\tau}{2\lVert s^{\prime}(x_{k})\nabla f(x_{k})\rVert}\,$,
where first inequality is due to $(1-\tau/\tau_{\mathcal{M}})^{-1}\leq 2$ and the second inequality due to Lemma 21. For brevity let us denote $y_{k}=x_{k}-\gamma_{k}s^{\prime}(x_{k})\nabla f(x_{k})$ and $z_{k}=x_{k}-\gamma_{k}P_{0}(x_{k})\nabla f(\pi(x_{k}))$. Also let $L_{0}=\operatorname{Lip}(\nabla(f\circ\pi)|_{\mathcal{T}(\tau)})$. By Lemma 10 and $(1-\tau/\tau_{\mathcal{M}})^{-1}\leq 2$ it follows that
@@ F.2 Proof of Theorem 5
F.3 Auxiliary results
F.3.1 A lemma on norms of orthogonal vectors
@@
F.3.2 Gaussian tail bounds
@@
Appendix G Further numerical experiments and implementation details
G.1 Optimization over $O(n)$
G.1.1 Implementation details
@@
G.1.2 Experiments
Refer to caption Refer to caption
Figure 3: Objective value vs. flow time $t$ for the orthogonal manifold for $n=10$ (left, different $\sigma>0$) and for $n=20$ (right, $\sigma=0.05$)
G.2 Data-driven reference tracking
G.2.1 Benchmark systems and tracking goals
@@ G.2.2 Implementation details
Diffusion and training parameters: Same as in Appendix G.1.1 with this time $N_{\operatorname{epochs}}=50000$ training epochs. For the DRGD step-size we pick a fixed step-size of $\gamma=0.001$.
G.2.3 Experiments

### v2

@@ Appendix G Discretized Riemannian gradient flow and descent
$\sup_{x\in\mathcal{T}(\tau)}\lVert\texttt{v}^{\prime}(x)\rVert_{\infty}\leq\epsilon+\frac{1}{1-\tau/\tau_{\mathcal{M}}}\,.$
@@
$\lVert G_{\texttt{v}}^{\eta}(x)\rVert\leq C\lVert\texttt{v}^{\prime}(x)\rVert+\eta\lVert\texttt{v}(x)-x\rVert\leq C(\epsilon+\frac{1}{1-\tau/\tau_{\mathcal{M}}})+2\eta\tau\text{\ \ for\ \ }x\in\mathcal{T}(\tau)$
@@
G.1 Discretized approximate Riemannian gradient flow
@@
G.2 Proof of Theorem 5
Let us write $\epsilon^{\prime}=\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3}\leq\tau/2$. First we note that $\{x_{k}\}_{k=0}^{\infty}\subseteq\mathcal{T}(\epsilon^{\prime})$ as soon as $x_{0}\in\mathcal{T}(\tau/2)$, because $\texttt{v}(\mathcal{T}(\tau))\subseteq\mathcal{T}(\epsilon^{\prime})$ and $x_{k}-\gamma_{k}\texttt{v}^{\prime}(x_{k})\nabla f(x_{k})\in\mathcal{T}(\tau)$, since
$\gamma_{k}\leq\frac{\tau}{2C(\epsilon^{\prime}+(1-\tau/\tau_{\mathcal{M}})^{-1})}\leq\frac{\tau}{2\lVert\texttt{v}^{\prime}(x_{k})\nabla f(x_{k})\rVert}\,$,
where first inequality is due to $(1-\tau/\tau_{\mathcal{M}})^{-1}\leq 2$ and the second inequality due to Lemma 23. For brevity let us denote $y_{k}=x_{k}-\gamma_{k}\texttt{v}^{\prime}(x_{k})\nabla f(x_{k})$ and $z_{k}=x_{k}-\gamma_{k}P_{0}(x_{k})\nabla f(\pi(x_{k}))$. Also let $L_{0}=\operatorname{Lip}(\nabla(f\circ\pi)|_{\mathcal{T}(\tau)})$. By Lemma 12 and $(1-\tau/\tau_{\mathcal{M}})^{-1}\leq 2$ it follows that
@@ G.2 Proof of Theorem 5
G.3 Auxiliary results
G.3.1 A lemma on norms of orthogonal vectors
@@
G.3.2 Gaussian tail bounds
@@
Appendix H Further numerical experiments and implementation details
H.1 Optimization over $O(n)$
H.1.1 Implementation details
@@
H.2 Data-driven reference tracking
H.2.1 Benchmark systems and tracking goals
@@ H.2.2 Implementation details
Diffusion and training parameters: Same as in Appendix H.1.1 with this time $N_{\operatorname{epochs}}=50000$ training epochs. For the DRGD step-size we pick a fixed step-size of $\gamma=0.001$.
H.2.3 Experiments